OpenStudy (anonymous):

can someone please help me to understand this?Which of the following shows the square completed in this equation? x2 + 3x - 5 = 0 (x + 9)2 = 14 (x - 3)2 = 8 (x + 1.5)2 = 7.25

OpenStudy (anonymous):

@amistre64 @mathstudent55

OpenStudy (flacalol):

C = (x + 1.5)2 = 7.25

OpenStudy (flacalol):

5. (X+1.5)^2=7.25 does. First, from X^2+3x-5=0, go to X^2+3x=5. Add (1.5)^2 or 2.25 to both sides and you would get (X+1.5)^2=7.25